What volume of a 0.169 M barium hydroxide solution is required to neutralize 15.2 mL of a 0.397 M hydrochloric acid solution?


_____________mL barium hydroxide

1
Expert's answer
2021-10-13T04:46:42-0400

Ba(OH)2 + 2HCl → BaCl2 + 2H2O

Proportion 1:

0.397 mol – 1000 mL

x mol – 15.2 mL

"x = \frac{0.397 \times 15.2}{1000} = 0.00603\;mol \;(HCl)"

According to the reaction equation:

n(Ba(OH)2) "= \frac{1}{2}n(HCl) = \frac{1}{2} \times 0.00603 = 0.00301 \;mol"

Proportion 2:

0.169 mol – 1000 mL

0.00301 – y mL

"y = \frac{0.00301 \times 1000}{0.169} = 17.853 \;mL"

Answer: 17.85 mL